# A flexible framework for structural plasticity in GPU-accelerated sparse spiking neural networks

### v1

@@ 3.2 Topographic map formation
We then measure the computational speed of the implementation. The runtime for increasing network sizes is illustrated in figure 9. At the network size used in the original publication [22] ($s=1$) we observe faster-than-realtime performance with the GeNN implementation. Even though the measured wall-clock time increases when increasing the network size, the proportionality factor is smaller than one: For example, at $s=7$ the network size grows by a factor of $49$, but the simulation time only by ca. $10$. The largest contributions to the total runtime are in fact due to the connectivity updates, in particular the time for remapping the inverse ragged matrix, followed by the update on the device required for STDP, while the time for the host update is negligible. The number of successful rewiring attempts remains constant when increasing the network size because the formation rule only establishes local connections. In contrast, the total number of rewiring attempts increases with the network size but is performed in parallel. Neuron and synapse updates (i.e., presynaptic and postsynaptic, including synaptic plasticity), and the Python code to switch the stimulus center location of the correlated input are also comparatively insignificant for larger network sizes. The gap between the sum of individual contributions and the total simulation time measured on the Python level is most likely due to the overheads of Python calling C++.
@@ 4 Discussion
We propose a flexible and efficient structural plasticity framework for the GeNN simulator which addresses the challenges of parallelizing structural plasticity algorithms and handles their complex dependencies on other neuronal and synaptic mechanisms, while minimizing the computational cost of changing connectivity in allocated data structures. We demonstrate this framework using two different applications: In a spike-based ML context, we train a recurrent classifier with very sparse connectivity using supervised learning and the DEEP R structural plasticity rule. In a computational neuroscience context, we show unsupervised learning leading to the refinement of a topographic map using a structural plasticity rule with activity-dependent and activity-independent contributions.
@@
The ‘lottery ticket hypothesis’[39] states that, within dense networks, there exist sparse sub-networks (the winning tickets) which can be trained in isolation to obtain the same test accuracy as the dense network in the same number of epochs. On this basis, the sparse classifiers we trained using DEEP R appear to be promising winning tickets as they maintain the same performance as networks with dense connectivity, even though they are extremely sparse: each hidden neuron only has around $5$ outgoing recurrent synapses and each input only around $25$ feed-forward synapses. However, as figure 6A shows, these potential winning tickets converge more slowly than the dense networks. This reduced convergence speed could be due to the ongoing rewiring process and it would be interesting to explore this further by retaining the connectivity structure of the sparse networks trained using DEEP R, reinitializing their weights and retraining them from scratch while keeping the connectivity static. [40] found that forming new connections randomly – as DEEP R does – limits the performance of sparse networks and propose using gradients from potential synapses to guide where new synapses are formed. While this approach is impossible in a truly sparse framework such as we are developing here, it would be interesting to combine DEEP R with more biologically-inspired ‘guidance’ such as the distance-dependant formation rule used in the topographic map formation model (table 1). Additionally, exploring how different tasks affect the efficacy of DEEP R and whether it can be combined with more efficient learning rules such as Eventprop [32, 16] would both be interesting future directions for this work.
@@
Using the novel structural plasticity framework, we have implemented a variation of the structural plasticity mechanism proposed by [22] which includes an interplay between structural plasticity, synaptic plasticity and spatial inter-neuron distance dependence. The mechanism promotes a refinement of receptive fields in a topographic map model [22] compatible with the results of [7]. Our analysis reveals further insights into the time course of synapse formation and elimination, in particular the feasibility of compensating for not fixing the maximum in-degree by adjusting the elimination probabilities. While the SpiNNaker implementation [7] achieves realtime performance with a simulation time-step of $h=1\,\mathrm{ms}$, GeNN demonstrates faster-than-realtime performance despite using a $10\times$ smaller time step. In contrast to previous implementation, the GeNN model applies multiple rewiring attempts in parallel, and demonstrates the scaling performance on a GPU upon upscaling the network size. At the original size, $10$ rewiring attempts are performed in parallel and at the largest tested size $10\times 49=490$. Other synaptic plasticity rules that involve more inherent parallelism would benefit even more from this feature. Both the structural plasticity mechanism and the network model used to demonstrate topographic map refinement serve as proofs-of-concept for the technological capability rather than claiming full biological realism. [22] and [7] provide extensive discussions on the limitations and on potential improvements, for example, axon branching could be guided by the existence of axons, homeostatic mechanisms could avoid the need for new synapses to start strong, and more realistic inputs, distance-dependent delays and inhibition could be added.
@@
The performance impact of structural plasticity updates on the two models we simulate is drastically different. In our e-prop classifier models, DEEP R only adds a $0.5\,\mathrm{\%}$ overhead to training time whereas, when simulating large topographic map models (scale $s=7$), structural plasticity updates account for around $90\,\mathrm{\%}$ of the simulation time. There are numerous reasons for this. Firstly, DEEP R is only run after each batch of $18456$ timesteps whereas the refinement rule used in the topographic map model is run much more frequently (every $10$ timesteps). Additionally, in each of these batches, $512$ model instances are updated in parallel further amortizing the cost of the DEEP R updates. The final issue is the large ‘remap time’ seen in figure 9 (see section 2.1). Because, only around $50\,\mathrm{\%}$ of connectivity updates in the topographic map model result in connection changes, one easy optimisation would be to only perform the remapping process if the connectivity has actually changed. Furthermore, more efficient GPU STDP implementations have been recently developed [41] which do not require this type of data structure at all. However, incorporating them into GeNN is not currently a priority and definitely beyond the scope of this paper.
@@
This study demonstrates the feasibility of expressing different structural plasticity rules [22, 20] using the syntax of our novel structural plasticity framework and enabling efficient GPU-based simulations of network dynamics. We hope that the proposed framework serves as a useful tool to support further explorations of biologically founded structural plasticity mechanisms and their exploitation for spike-based ML scenarios. Despite considering generic challenges of structural plasticity implementations, our design is tailored towards GPU-based systems and in particular the GeNN simulator. The suggested approach of identifying the simulator-specific demands and designing suitable structural plasticity functionality, however, is of a more general nature, and we are curious how it can be adapted to other simulation software and neuromorphic hardware platforms.

### v2

@@ 3.2 Topographic map formation
We then measure the computational speed of the implementation. The runtime for increasing network sizes is illustrated in figure 10. At the network size used in the original publication [undefaa] ($s=1$) we observe faster-than-realtime performance with the GeNN implementation. Even though the measured wall-clock time increases when increasing the network size, the proportionality factor is smaller than one: For example, at $s=7$ the network size grows by a factor of $49$, but the simulation time only by ca. $10$. The largest contributions to the total runtime are in fact due to the connectivity updates, in particular the time for remapping the inverse ragged matrix, followed by the update on the device required for STDP, while the time for the host update is negligible. The number of successful rewiring attempts remains constant when increasing the network size because the formation rule only establishes local connections. In contrast, the total number of rewiring attempts increases with the network size but is performed in parallel. Neuron and synapse updates (i.e., presynaptic and postsynaptic, including synaptic plasticity), and the Python code to switch the stimulus center location of the correlated input are also comparatively insignificant for larger network sizes. The gap between the sum of individual contributions and the total simulation time measured on the Python level is most likely due to the overheads of Python calling C++.
@@ 4 Discussion
We propose a flexible and efficient structural plasticity framework for the GeNN simulator which addresses the challenges of parallelizing structural plasticity algorithms and handles their complex dependencies on other neuronal and synaptic mechanisms, while minimizing the computational cost of changing connectivity in allocated data structures. We demonstrate this framework using two different applications: In a spike-based ML context, we train recurrent classifiers with very sparse connectivity using supervised learning and the DEEP R structural plasticity rule. In a computational neuroscience context, we show unsupervised learning leading to the refinement of a topographic map using a structural plasticity rule with activity-dependent and activity-independent contributions.
@@
The ‘lottery ticket hypothesis’[undefar] states that, within dense networks, there exist sparse sub-networks (the winning tickets) which can be trained in isolation to obtain the same test accuracy as the dense network in the same number of epochs. On this basis, the sparse classifiers we trained using DEEP R appear to be promising winning tickets as they maintain the same performance as networks with dense connectivity, even though they are extremely sparse: each hidden neuron only has around $5$ outgoing recurrent synapses and each input only around $25$ feed-forward synapses. However, as figure 6A shows, these potential winning tickets converge more slowly than the dense networks. This reduced convergence speed could be due to ‘competition’ between e-prop learning, L1 regularization and the rewiring process (which occurs every batch) and it would be interesting to explore this further by retaining the connectivity structure of the sparse networks trained using DEEP R, reinitializing their weights and retraining them from scratch while keeping the connectivity static. [undefas] found that forming new connections randomly – as DEEP R does – limits the performance of sparse networks and propose using gradients from potential synapses to guide where new synapses are formed. While this approach is impossible in a truly sparse framework such as we are developing here, it would be interesting to combine DEEP R with more biologically-inspired ‘guidance’ such as the distance-dependent formation rule used in the topographic map formation model (table 1). Additionally, exploring how different tasks affect the efficacy of DEEP R and whether it can be combined with more efficient learning rules such as Eventprop [undefak, undeft] would both be interesting future directions for this work.
@@
Using the novel structural plasticity framework, we have implemented a variation of the structural plasticity mechanism proposed by [undefaa] which includes an interplay between structural plasticity, synaptic plasticity and spatial inter-neuron distance dependence. The mechanism promotes a refinement of receptive fields in a topographic map model [undefaa] compatible with the results of [undefj]. Our analysis reveals further insights into the time course of synapse formation and elimination, in particular the feasibility of compensating for not fixing the maximum in-degree by adjusting the elimination probabilities. While the SpiNNaker implementation [undefj] achieves realtime performance with a simulation time-step of $h=1\,\mathrm{ms}$, GeNN demonstrates faster-than-realtime performance despite using a $10\times$ smaller time step. In contrast to previous implementations, the GeNN model applies multiple rewiring attempts in parallel, and demonstrates the scaling performance on a GPU upon upscaling the network size. At the original size, $10$ rewiring attempts are performed in parallel and at the largest tested size $10\times 49=490$. Other synaptic plasticity rules that involve more inherent parallelism would benefit even more from this feature. Both the structural plasticity mechanism and the network model used to demonstrate topographic map refinement serve as proofs-of-concept for the technological capability rather than claiming full biological realism. [undefaa] and [undefj] provide extensive discussions on the limitations and on potential improvements, for example, axon branching could be guided by the existence of axons, homeostatic mechanisms could avoid the need for new synapses to start strong, and more realistic inputs, distance-dependent delays and inhibition could be added. Any biological interpretation of our example results should take the simplifications of the current approach into account.
@@
The performance impact of structural plasticity updates on the two models we simulate is drastically different. In our e-prop classifier models, DEEP R only adds a $0.5\,\mathrm{\%}$ overhead to training time whereas, when simulating large topographic map models (scale $s=7$), structural plasticity updates account for around $90\,\mathrm{\%}$ of the simulation time. There are numerous reasons for this. Firstly, DEEP R is only run after each batch ($18456$ timesteps for DVS-gesture) whereas the refinement rule used in the topographic map model is run much more frequently (every $10$ timesteps). Additionally, in each of these batches, $512$ model instances are updated in parallel further amortizing the cost of the DEEP R updates. The final issue is the large ‘remap time’ seen in figure 10 (see section 2.1). Because, only around $50\,\mathrm{\%}$ of connectivity updates in the topographic map model result in connection changes, one easy optimisation would be to only perform the remapping process if the connectivity has actually changed. Furthermore, more efficient GPU STDP implementations have been recently developed [undefat] which do not require this type of data structure at all. However, incorporating further optimisations into GeNN is beyond the scope of this paper and, while these overheads may limit the practicability of our approach for large models using STDP with a postsynaptic spike-triggered term, the performance impact of structural plasticity will be minimal for many other models.
@@
This study demonstrates the feasibility of expressing different structural plasticity rules [undefaa, undefx] using the syntax of our novel structural plasticity framework and enabling efficient GPU-based simulations of network dynamics. We hope that the proposed framework serves as a useful tool to support further explorations of biologically founded structural plasticity mechanisms and their exploitation for spike-based ML. While we have not explored it here, this framework could also be used to implement neurogenesis by instantiating the final number of neurons at the start of the simulation without any synaptic connections and connecting them over time. This approach has been widely used to implement neurogenesis both in software simulators [undefau] and neuromorphic hardware [undefav].
Despite considering generic challenges of structural plasticity implementations, our design is tailored towards GPU-based systems and in particular the GeNN simulator. However, such frameworks are potentially applicable to other simulation software that uses code generation such as Brian 2 [undefaw] or NESTML [undefax] as well as programmable neuromorphic hardware. For example, ongoing work [undefay] to target the SpiNNaker [undefp] and SpiNNaker 2 [undefaz] neuromorphic systems using NESTML would provide an interesting route for running flexible structural plasticity rules at scale. Furthermore, both Loihi [undefaaa] and BrainScaleS–2 [undefq] supports sparse connectivity and, potentially, code could be generated to run on the general processors present in both systems to modify this connectivity. Accelerated-time systems such as BrainScaleS–2 would be a particularly interesting choice for implementing structural plasticity as they can overcome the long simulation times otherwise required to simulate the slow speed of structural changes in biology.